How doe subnetting segments a network to enable better communication.

How do you calculate IPv4 subnets for a /24 prefix.

How do you calculate IPv4 subnets for a /16 and /8 prefix..

How do you create a flexible addressing scheme using variable length subnet masking (VLSM).
Answer #1

Subnetting is basically just a way of splitting a TCP/IP network into smaller, more manageable pieces. The basic idea is that if you have an excessive amount of traffic flowing across your network, then that traffic can cause your network to run slowly. When you subnet your network, you are splitting the network into a separate, but interconnected network. That way, most of the network traffic will be isolated to the subnet in which it originated. Of course you can still communicate across a subnet, but the only time that traffic will cross subnet boundaries is when it is specifically destined for a host residing in an alternate subnet.

Here are five subnetting benefits you should consider

1. Improve network performance and speed

A single broadcast packet sends out information that reaches every device connected to that network because each device has an entry point into the network. A large number of entry points, however, can negatively impact internetwork switching device performance, as well as your network’s overall performance.

Another issue with broadcast packets is that they can spam every device within a network, even devices that aren’t relevant to the task at hand, which can strain a network’s capacity and cause it to collapse.

But subnetting enables you to ensure that information remains in the subnetted network or broadcast domain, which allows other subnets to maximize their speed and effectiveness. Subnetting also divides your network’s broadcast domains, enabling you to better control traffic flow, thus increasing network performance!

A word of caution, though. You’re better off limiting traffic to a single subnet instead of letting it move from subnet to subnet. So, you should limit the number of devices on your subnet whenever possible, along with controlling the traffic flow between subnets. Doing this will improve your network’s speed and performance.

2. Reduce network congestion

Subnetting ensures that traffic destined for a device within a subnet stays in that subnet, which reduces congestion. Through strategic placement of subnets, you can help reduce your network’s load and more efficiently route traffic.

So, what happens to a large network with no subnets? Every computer would see broadcast packets from all the computers and servers on the network, resulting in the switches having to move all that traffic to the appropriate ports. This leads to increased congestion, reduced network performance, and slower response times.

However, using a router to move traffic between subnets results in no broadcast traffic or any information that doesn’t need to be routed being moved to other subnets. Because the amount of traffic within each subnet is reduced, the speed of each subnet is increased, which eases network congestion.

3. Boost network security

You might be thinking, “What if a device in my network is comprised?” By splitting your network into subnets, you can control the flow of traffic using ACLs, QoS, or route-maps, enabling you to identify threats, close points of entry, and target your responses more easily.

You also can split your network using routers to connect subnets though the configuration of ACLs on the routers and switches. As a result, devices in a subnet are unable to access the entire network.

Another option is to limit access to resources on wireless clients, ensuring that valuable information isn’t easily accessible in remote locations.

4. Control network growth

When you're planning and designing a network, size is something that needs to be taken into consideration. One of the key benefits of subnetting is that it enables you to control the growth of your network.

You can use a popular host formula to determine the size of your network. Take the number of zeros in the mask of your subnet when converted to binary, take two to the power of that number, then minus two — and then you will have the number of possible hosts for that subnet mask. That was a bit of a doozy, so here’s a more in-depth explanation of the host formula.

Your next step is to figure out the expected growth of the network, which in most cases will be based heavily on the physical size of your building. For example, will the number of devices needed remain steady or could it eventually double? If so, you will need to adjust the equation for the host formula accordingly, in order to determine the proper IP address space for your network.

5. Ease administration

Are you a network admin? Then subnetting is a no-brainer because it can make your job a lot easier. By subnetting, you can create networks that have more logical host limits, as opposed to the limitations of IP addressing classes: Eight bits for Class A, 16 bits for Class B, and 24 bits for Class C. Think about it this way, if the internet was limited to only those three classes, every network would have only 254, 64,000, or 16 million IP addresses for host devices.

In the absence of subnets, networks with more than 254 devices need a Class B allocation, which can waste thousands of IP addresses. By subnetting, you can select the number of bits in your subnetwork, creating more realistic host limits.

Subnetting also is an effective way to keep tabs on the machines on your network, which in turn can help you determine which machines need attention should problems arise. So while they take careful planning and implementation, subnetted networks generally are easier to manage and troubleshoot.

-----------------------------------------------------------------------------------------------------------------------------------------------------------------------

Prefix length /24 = 255.255.255.0
/24 = 32-24 = 8 host bit
2ˆ8 = 256
So in /24 has 256 IP, 1 broadcast, 1 network so usable is 254 IP

Number of Subnets in an Address Block

Given an address block (network/prefix length), you can determine the number of subnets that can be gotten from that address block as long as you know the subnet size requirements. The formula for this is:

For example, we can get sixteen (16) /28 subnets from a /24 reference address block:

List of Subnets in an Address block

we determined the number of subnets that can be gotten from a particular address block. Now, we need to determine what those subnets actually are. To do this, we need to know the following things:

  1. The octet in which a subnet exists
    • 1st octet: /1 to /8
    • 2nd octet: /9 to /16
    • 3rd octet: /17 to /24
    • 4th octet: /25 to /32
  2. The maximum number of bits in the boundary (octet) in which the subnet belongs
    • 1st octet: 8
    • 2nd octet: 16
    • 3rd octet: 24
    • 4th octet: 32
  3. The block size of the subnet

For example, a /28 subnet exists in the 4th octet. The maximum number of bits in that octet is 32. Therefore, the block size is:

Here’s another example. A /18 subnet exists in the 3rd octet. The maximum number of bits in that octet is 24. Therefore, the block size is:

We can now use this knowledge to list the subnets in a particular address block. For example, what are the /27 subnets that exist in the 174.53.4.0/24 address block?

First, we know that there are eight (8) /27 subnets in a /24 address block (i.e. 227-24 = 23 = 8). Next, we can see that the /27 subnet exists in the 4th octet. The maximum number of bits in that octet is 32. Therefore, the block size is:

Knowing this, we can now list the subnets by starting at first network of the given blockand incrementing by the block size in the 4th octet:

  1. 174.53.4.0/27
  2. 174.53.4.32/27
  3. 174.53.4.64/27
  4. 174.53.4.96/27
  5. 174.53.4.128/27
  6. 174.53.4.160/27
  7. 174.53.4.192/27
  8. 174.53.4.224/27

Let’s take another example. List the /23 subnets that exist in the 141.67.128.0/21 address block.

First, we know that there are four (4) /23 subnets in a /21 address block. Next, we can see that the /23 subnet exists in the 3rd octet. The maximum number of bits in that octet is 24. Therefore, the block size is:

Knowing this, we can now list the subnets by starting at first network of the given blockand incrementing by the block size in the 3rd octet:

  • 141.67.128.0/23
  • 141.67.130.0/23
  • 141.67.132.0/23
  • 141.67.134.0/23

Warning: You need to be careful about the “first network of the given block”. It must be a multiple of the block size. For example, 128 is a multiple of 2. If you are not sure, start at 0 and increase by the block size. For example, 141.67.0.0, 141.67.2.0, 141.67.4.0, …, 141.67.126.0, 141.67.128.0, and so on.

Exercise #1: List the /13 subnets that exist in the 131.80.0.0/12 address block. (The answer is at the end of the article.)

Address Range of Subnet

When you know the size of a subnet, it becomes easy to determine the valid addresses in that subnet. We just need to add one (1) IP address to the subnet address and subtract two (2) IP addresses from the next subnet address. We add 1 because the first address is the network address and we subtract 2 instead of 1 because the last address in a subnet is the broadcast address.

Note: The next subnet address is just the subnet plus the block size. Keep in mind that this “next subnet address” may not be a valid address block.

For example, what is the valid address range of the 141.67.132.0/23 above? Since the block size is 2, we know that the next subnet is 141.67.134.0/23. Therefore, the valid address range is:

  • Start address: 141.67.132.0 + 1 = 141.67.132.1
  • End address: 141.67.134.0 – 2 = 141.67.133.254
  • Broadcast address: 141.67.134.0 – 1 = 141.67.133.255

Here’s another example. What is the valid address range of the 10.192.0.0/10 subnet? Since the block size is 64, we know that the next subnet will be 10.256.0.0/10. Therefore, the valid address range is:

  • Start address: 10.192.0.0 + 1 = 10.192.0.1
  • End address: 10.256.0.0 – 2 = 10.255.255.254
  • Broadcast address: 10.256.0.0 – 1 = 10.255.255.255

-----------------------------------------------------------------------------------------------------------------------------------------------------------------------

Prefix length /16 = 255.255.0.0
/16 = 32-16 = 16 host bit
2ˆ16 = 65536
So in /16 has 65536 IP, 1 broadcast, 1 network so usable is 65534 IP

Number of Subnets in an Address Block

Given an address block (network/prefix length), you can determine the number of subnets that can be gotten from that address block as long as you know the subnet size requirements. The formula for this is:

For example, we can get sixteen (256) /24 subnets from a /16 reference address block:

Number of /24 subnets from /16 block = 2 ^24-16 = 256

Prefix length /8 = 255.0.0.0
/8 = 32-8 = 24 host bit
2ˆ24 = 16777216
So in /8 is 16777216 IP, 1 broadcast, 1 network so usable is 16777214 IP

Number of Subnets in an Address Block

Given an address block (network/prefix length), you can determine the number of subnets that can be gotten from that address block as long as you know the subnet size requirements. The formula for this is:

For example, we can get sixteen (65536) /24 subnets from a /16 reference address block:

Number of /24 subnets from /8 block = 2 ^24-8 = 65536

-----------------------------------------------------------------------------------------------------------------------------------------------------------------------

VLSM Addressing Schemes

Variable-length subnet masking (VLSM) subnetting is similar to traditional subnetting in that bits are borrowed to create subnets. The formulas to calculate the number of hosts per subnet, and the number of subnets created still apply. The difference is that subnetting is not a single-pass activity.

VLSM Review

You probably noticed that the starting address space in Subnetting Scenario 3 is not an entire classful address. In fact, it is subnet 5 from Subnetting Scenario 2. So in Subnetting Scenario 3, you “subnetted a subnet.” That is what VLSM is in a nutshell: subnetting a subnet.

Let’s use a small example. Given the address space 172.30.4.0/22 and the network requirements shown in Figure 9-1, apply an addressing scheme that conserves the most amount of addresses for future growth.

Figure 9-1 VLSM Example Topology

We need five subnets: four LAN subnets and one WAN subnet. Starting with the largest host requirement on LAN 3, begin subnetting the address space.

To satisfy the 250 hosts requirement, we leave 8 hosts bits (28 – 2 = 254 hosts per subnet). Because we have 10 host bits total, we borrow 2 bits to create the first round of subnets (22= 4 subnets). The starting subnet mask is /22 or 255.255.252.0. We turn on the next two bits in the subnet mask to get /24 or 255.255.255.0. The multiplier is 1. The four subnets are as follows:

  • Subnet 0: 172.30.4.0/24
  • Subnet 1: 172.30.5.0/24
  • Subnet 2: 172.30.6.0/24
  • Subnet 3: 172.30.7.0/24

Assigning Subnet 0 to LAN 3, we are left with three /24 subnets. Continuing on to the next largest host requirement on LAN 4, we take Subnet 1, 172.30.5.0/24, and subnet it further.

To satisfy the 100 hosts requirement, we leave 7 bits (27 – 2 = 128 hosts per subnet). Because we have 8 host bits total, we can borrow only 1 bit to create the subnets (21 = 2 subnets). The starting subnet mask is /24 or 255.255.255.0. We turn on the next bit in the subnet mask to get /25 or 255.255.255.128. The multiplier is 128. The two subnets are as follows:

  • Subnet 0: 172.30.5.0/25
  • Subnet 1: 172.30.5.128/25

Assigning Subnet 0 to LAN 4, we are left with one /25 subnet and two /24 subnets. Continuing on to the next largest host requirement on LAN 1, we take Subnet 1, 172.30.5.128/25, and subnet it further.

To satisfy the 60 hosts requirement, we leave 6 bits (26 – 2 = 62 hosts per subnet). Because we have 7 host bits total, we borrow 1 bit to create the subnets (21 = 2 subnets). The starting subnet mask is /25 or 255.255.255.128. We turn on the next bit in the subnet mask to get /26 or 255.255.255.192. The multiplier is 64. The two subnets are as follows:

  • Subnet 0: 172.30.5.128/26
  • Subnet 1: 172.30.5.192/26

Assigning Subnet 0 to LAN 1, we are left with one /26 subnet and two /24 subnets. Finishing our LAN subnetting with LAN 2, we take Subnet 1, 172.30.5.192/26, and subnet it further.

To satisfy the 10 hosts requirement, we leave 4 bits (24 – 2 = 14 hosts per subnet). Because we have 6 host bits total, we borrow 2 bits to create the subnets (22 = 4 subnets). The starting subnet mask is /26 or 255.255.255.192. We turn on the next two bits in the subnet mask to get /28 or 255.255.255.240. The multiplier is 16. The four subnets are as follows:

  • Subnet 0: 172.30.5.192/28
  • Subnet 1: 172.30.5.208/28
  • Subnet 2: 172.30.5.224/28
  • Subnet 3: 172.30.5.240/28

Assigning Subnet 0 to LAN 2, we are left with three /28 subnets and two /24 subnets. To finalize our addressing scheme, we need to create a subnet only for the WAN link, which needs only two host addresses. We take Subnet 1, 172.30.5.208/28, and subnet it further.

To satisfy the two hosts requirement, we leave 2 bits (22 – 2 = 2 hosts per subnet). Because we have 4 host bits total, we borrow 2 bits to create the subnets (22 = 4 subnets). The starting subnet mask is /28 or 255.255.255.240. We turn on the next 2 bits in the subnet mask to get /30 or 255.255.255.252. The multiplier is 4. The four subnets are as follows:

  • Subnet 0: 172.30.5.208/30
  • Subnet 1: 172.30.5.212/30
  • Subnet 2: 172.30.5.216/30
  • Subnet 3: 172.30.5.220/30

We assign Subnet 0 to the WAN link. We are left with three /30 subnets, two /28 subnets, and two /24 subnets.